Explain what is meant by (1) a period-by-period return and (2) a multi-period return.

Explain (1) the difference between time-weighted and dollar- or money-weighted and (2) how do these equate to the two returns in Question 1?
Answer #1

A period by period return is computed for each period.

For example if in period 0, there is a cash inflow of 500

and at period 1 this becomes 750, then period return becomes (750-500)/500= 50%

In period 2, if the value of investment becomes 700, then second period return is -50/750= -6.67%

Multiperiod return is (700/500)^0.5-1= 18.32%

The time weighted dollar return is the compound annual growth rate of the portfolio and is similar to the multiperiod return shown above.

The dollar weighted return takes into account the cash inflows and outflows and is akin to the internal rate of return (IRR) the cash flows. If we adjust any cash inflow or outflow in the period by period return, then IRR is similar to the period by period return